# Stability via resampling: statistical problems beyond the real line

### v1

@@ 1.1 Problem settings
In Bayesian data analysis, given a data set $\mathcal{D}$ containing $n$ observations drawn from some distribution $p_{\theta}$, we are interested in the posterior distribution of the parameter $\theta$. For instance, $\theta$ might represent a particular model from a finite set of possibilities, in which case the posterior distribution on $\theta$ is simply a vector of weights on the finite set of possibilities. Alternatively, $\theta$ might take values in $\mathbb{R}^{d}$ (e.g., coefficients in some regression model), in which case the posterior distribution is a distribution over $\mathbb{R}^{d}$—a more complex object.
@@ 1.2 Our contributions
Density estimation $L$-Lipschitz densities on $[0,1]$ Total variation dist. $\mathsf{d}_{\mathsf{TV}}$
@@
Table 1: Example problem settings for our stability framework. The first three examples correspond to Problem Settings 1-3 described in Section 1.
@@ 2.1 Defining stability
where $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ is the output of the algorithm run on the full data set $\mathcal{D}$, $\hat{w}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$ is the output of the algorithm run on the leave-one-out data set $\mathcal{D}^{\setminus i}=(Z_{j})_{j\neq i}$ and the expectation averages over $\xi\sim\text{Uniform}[0,1]$.
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}(\hat{w},\hat{w}^{\setminus i})>\varepsilon\right\}\leq\delta,$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}=\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon^{2}\right\}\right]\\ \leq\varepsilon^{-2}\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\varepsilon^{-2}\beta^{2}.$
For the second claim, since $\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\leq\varepsilon^{2}+\mathbf{1}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$, we have
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi}\,\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)$ $\leq\varepsilon^{2}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)>\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$
@@ 3 Constructing a bagged algorithm via resampling
Define the set of finite sequences of indices in $[n]$:
@@ 3.1 A general framework for bagging
The four conditions of 2 are all satisfied by the standard resampling schemes described above, with $p_{\mathcal{Q}_{n}}=1-\left(1-\frac{1}{n}\right)^{m}$ for bagging (sampling $m$ out of $n$ points uniformly with replacement), and $p_{\mathcal{Q}_{n}}=\frac{m}{n}$ for subbagging (sampling $m$ out of $n$ points uniformly without replacement).
@@ 4 Main results for a Hilbert space
  • If $\mathcal{W}\subset L_{2}(\mathbb{R}^{d})$, the space of square-integrable functions $\mathbb{R}^{d}\to\mathbb{R}$, then the Hilbert norm is given by the $L_{2}$ norm on functions, $\|w\|_{\mathcal{H}}=(\int_{\mathbb{R}^{d}}w(x)^{2}\;\mathsf{d}x)^{1/2}$. Alternatively, we can take a Sobolev space of functions $\mathbb{R}^{d}\to\mathbb{R}$ with smoothness parameter $s$ with the Hilbert norm being the Sobolev norm $\|\cdot\|_{s,2}$ in (3); see Experiment 3 in Section 6.1 for details.
@@
Corollary 7 generalizes the main result of Soloff et al. (2023), which focuses on the supervised learning setting where the data points $Z_{i}=(X_{i},Y_{i})$ contain features $X_{i}\in\mathcal{X}$ and a response $Y_{i}\in\mathbb{R}$, and the output of the algorithm $\mathcal{A}$ is a predicted value $\hat{f}(x)$, at a test point $x\in\mathcal{X}$—i.e., in that work, the output space $\mathcal{H}=\mathbb{R}$.
@@
The main result of (Soloff et al., 2023, Theorem 8), ensuring the stability of algorithms that return a real-valued output, can be recovered as a special case of Corollary 7 above, by taking $\mathcal{W}=[0,1]\subseteq\mathcal{H}=\mathbb{R}$ (note that we then have $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})=\frac{1}{2}$). But remarkably, the result of Theorem 6 shows that there is no price to pay for high-dimensionality (or even for infinite dimensionality): the stability guarantee is exactly the same for any Hilbert space, regardless of dimension, as long as the radius of the output space $\mathcal{W}$ is bounded.
@@ 4.1 Results for a finite number of bags
The results above apply to the derandomized bagged version of $\mathcal{A}$, as constructed in Definition 5. In practice, of course, it is not feasible to calculate this output $\hat{w}$ since we would need to average over all possible bags $r$, so we would instead use the bagged estimator $\hat{w}$ constructed in Definition 4 for some finite number of bags $B$. Since we are simply taking a Monte Carlo approximation to the derandomized bagged estimator of Definition 5, we would expect to incur error on the order of $\frac{1}{\sqrt{B}}$. To formalize this, we first state a generalization of the Azuma-Hoeffding inequality:
@@
Let $\hat{w}_{1},\ldots,\hat{w}_{B}\in\mathcal{H}$ be independent random variables in a Hilbert space $\mathcal{H}$ such that $\|\hat{w}_{b}-\mathbb{E}\hat{w}_{b}\|_{\mathcal{H}}\leq C$ almost surely for $b=1,\ldots,B$. Then for any $\delta>0$,
$\mathbb{P}\left\{\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}\hat{w}_{b})\bigg{\|}_{\mathcal{H}}\leq C\sqrt{\frac{1}{2B}\log\left(\frac{2e^{2}}{\delta}\right)}\right\}\geq 1-\delta.$
@@
$\mathbb{E}\left[\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}\hat{w}_{b})\bigg{\|}_{\mathcal{H}}^{2}\right]\leq\frac{C^{2}e^{2}}{B}.$
@@ 5.1 Theoretical guarantee
For instance, returning to our earlier example, if $\mathcal{W}=\Delta_{K-1}$ (and the norm is given by $\mathsf{d}_{\mathsf{TV}}$) then we can take $v_{j}$ to be the $j$th canonical basis vector, and then 2 holds, with this same value of $K$, and with $R=1$ and $\rho=0$. Here $R$ is playing a role analogous to the radius, $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})$, in the Hilbert space setting.
@@ 5.2 Examples
Next, consider a continuous setting, where $\mathcal{A}$ returns a density on $[0,1]$, given by some function $\hat{w}:[0,1]\rightarrow[0,\infty)$ satisfying $\int_{t=0}^{1}\hat{w}(t)\;\mathsf{d}t=1$. In this case we have $\mathsf{d}_{\mathsf{TV}}(w,w^{\prime})=\frac{1}{2}\int_{t=0}^{1}|w(t)-w^{\prime}(t)|\;\mathsf{d}t$.
@@
$\mathcal{B}=L_{1}([0,1]):=\left\{f:[0,1]\to\mathbb{R}:\int_{t=0}^{1}|f(t)|\;\mathsf{d}t<\infty\right\},$
equipped with the norm $\|f\|=\frac{1}{2}\int_{t=0}^{1}|f(t)|\;\mathsf{d}t$ (note that this differs from the usual $\ell_{1}$ norm by the factor $\frac{1}{2}$, in order to agree with total variation distance). Suppose that our algorithm returns densities that are further constrained to be $L$-Lipschitz, so that we have
@@
As for Problem Setting 2(a), we again begin by asking whether the Hilbert space result of Theorem 6 might already be sufficient in this setting. Indeed, while $\mathcal{B}=L_{1}([0,1])$ is not a Hilbert space, the Lipschitz constraint on $\mathcal{W}$ means that $\int_{t=0}^{1}|w(t)|^{2}\;\mathsf{d}t<\infty$ for all $w\in\mathcal{W}$, meaning that $\mathcal{W}$ can be viewed as a subset of the Hilbert space $L_{2}([0,1])$, equipped with norm $\|f\|_{\mathcal{H}}=\|f\|_{L_{2}}=(\int_{t=0}^{1}|f(t)|^{2}\;\mathsf{d}t)^{1/2}$. A straightforward calculation verifies that $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})\propto L^{1/4}$, meaning that by Theorem 6, derandomized bagging satisfies mean-square stability with respect to $\|\cdot\|_{L_{2}}$ for $\beta^{2}\propto L^{1/2}/n$. Since for this function space we have $\|\cdot\|_{L_{1}}\leq\|\cdot\|_{L_{2}}$, this immediately implies that mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ also holds with parameter $\beta^{2}\propto L^{1/2}/n$. However, if the Lipschitz constant $L$ is large, this may not be a satisfying result.
@@
Stability guarantee via
Theorem 6 (Hilbert space)
Stability guarantee via
Theorem 10 (Banach space)
@@ 6.1 Data and Methods
with $n=500$ and $d=40$. We then set $Y_{i}=\frac{\tilde{Y}_{i}-\min_{j}\tilde{Y}_{j}}{\max_{j}\tilde{Y}_{j}-\min_{j}\tilde{Y}_{j}}$ so that each response $Y_{i}$ is in the unit interval $[0,1]$. Note that the algorithm has access to the observed data $\mathcal{D}=(X_{i},Y_{i})_{i\in[n]}$, i.e., $\alpha_{i}$ and $\gamma_{i}$ are latent variables used only to generate the data $\mathcal{D}$. We apply sklearn.tree.DecisionTreeRegressor to train the regression trees, setting max_depth=50 and leaving all other parameters at their default values. We evaluate the stability of the learned function in the $L_{2}$ norm $\|\hat{w}\|_{L_{2}}=\int_{[0,1]^{d}}\hat{w}(x)\textnormal{d}x$.
@@
For this experiment, we reproduce Abadie et al. (2015)’s SCM analysis on the 1990 reunification of Germany, as discussed above in Section 1.1 (see Problem Setting 3). The data contains $n=16$ countries (aside from Germany), with $6$ predictors measured for each country (pre-treatment averages of various macroeconomic factors such as inflation and trade openness—see Table 1 of Abadie et al. (2015) for more details). SCM then estimates a weighted combination of these $n$ countries that represents a ‘synthetic West Germany’, meaning that the output of the algorithm is a vector $\hat{w}\in\Delta_{n-1}$. When we run SCM on a subset of the $n=16$ countries, we assign a weight of 00 to each omitted country. We run SCM using the Python package pysyncon and make extensive use of their reproduction of the Abadie et al. (2015) study. We evaluate the stability of these weights in the Euclidean norm $\|\cdot\|_{2}$.
@@
Figure 1: Simulation results comparing the stability of subbagging to that of the corresponding base algorithm under the given norm. Left: Histogram of leave-one-out perturbations. Right: for each $\varepsilon$, the smallest $\delta$ such that the algorithm has tail stability $(\varepsilon,\delta)$ in the sense of Definition 2. Higher curves thus represent greater instability. In the right panels we also plot the stability guarantee of Theorem 6 for Experiments 1-3 and Theorem 10 for Experiment 4 (the guarantees for derandomized bagging). For all four experiments, $m=n/2$ and $B=10,000$.

### v2

@@ 1.1 Problem settings
In Bayesian data analysis, given a data set $\mathcal{D}$ containing $n$ observations drawn from some distribution $p_{\theta}$, we are interested in the posterior distribution of the parameter $\theta$. For instance, $\theta$ might represent a particular model from a finite set of possibilities, in which case the posterior distribution on $\theta$ is simply a vector of weights on this finite set. Alternatively, $\theta$ might take values in a continuous space such as $[0,1]$ (e.g., if $\theta$ represents a probability) or $\mathbb{R}^{d}$ (e.g., coefficients for some regression model), in which case the posterior distribution is a distribution over $[0,1]$ or over $\mathbb{R}^{d}$—a more complex object.
@@ 1.2 Our contributions
Density estimation
(e.g., Bayesian posterior)
$L$-Lipschitz densities on $[0,1]$ Total variation dist. $\mathsf{d}_{\mathsf{TV}}$
@@
Table 1: Example problem settings for our stability framework, including the examples highlighted in Problem Settings 1–3 (in Section 1).
@@ 2.1 Defining stability
where $\hat{w}=\mathcal{A}(\mathcal{D};\xi)$ is the output of the algorithm run on the full data set $\mathcal{D}$, $\hat{w}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$ is the output of the algorithm run on the leave-one-out data set $\mathcal{D}^{\setminus i}=(Z_{j})_{j\neq i}$, and the expectation averages over $\xi\sim\text{Uniform}[0,1]$.
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}(\hat{w},\hat{w}^{\setminus i})\geq\varepsilon\right\}\leq\delta,$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon\right\}$ $=\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon^{2}\right\}\right]$
$\leq\varepsilon^{-2}\mathbb{E}_{\xi}\left[\frac{1}{n}\sum_{i=1}^{n}\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]\leq\varepsilon^{-2}\beta^{2}.$
For the second claim, since $\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\leq\varepsilon^{2}+\mathbf{1}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$, we have
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\xi}\left[\mathsf{d}^{2}\left(\hat{w},\hat{w}^{\setminus i}\right)\right]$ $\leq\varepsilon^{2}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{\xi}\left\{\mathsf{d}\left(\hat{w},\hat{w}^{\setminus i}\right)\geq\varepsilon\right\}\cdot\sup_{w,w^{\prime}\in\mathcal{W}}\mathsf{d}^{2}\left(w,w^{\prime}\right)$
@@ 3 Constructing a bagged algorithm via resampling
Following the framework and notation of Soloff et al. (2024), we define the set of finite sequences of indices in $[n]$:
@@ 3.1 A general framework for bagging
The four conditions of 2 are all satisfied by the standard resampling schemes described above, with $p_{\mathcal{Q}_{n}}=1-\left(1-\frac{1}{n}\right)^{m}$ for bootstrapping (sampling $m$ out of $n$ points uniformly with replacement), and $p_{\mathcal{Q}_{n}}=\frac{m}{n}$ for subbagging (sampling $m$ out of $n$ points uniformly without replacement).
@@ 4 Main results for a Hilbert space
  • If $\mathcal{W}\subset L_{2}(\mathbb{R}^{d})$, the space of square-integrable functions $\mathbb{R}^{d}\to\mathbb{R}$, then the Hilbert norm is given by the $L_{2}$ norm on functions, $\|w\|_{\mathcal{H}}=(\int_{\mathbb{R}^{d}}w(x)^{2}\;\textnormal{d}x)^{1/2}$. Alternatively, we can take a Sobolev space of functions $\mathbb{R}^{d}\to\mathbb{R}$ with smoothness parameter $s$ with the Hilbert norm being the Sobolev norm $\|\cdot\|_{s,2}$ (defined formally later on in (3)—see Experiment 3 in Section 6.1 for details).
@@
The main result of (Soloff et al., 2024, Theorem 8), ensuring the stability of algorithms that return a real-valued output, can be recovered as a special case of Corollary 7 above, by taking $\mathcal{W}=[0,1]\subseteq\mathcal{H}=\mathbb{R}$ (note that we then have $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})=\frac{1}{2}$). But remarkably, the result of Theorem 6 shows that there is no price to pay for high-dimensionality (or even for infinite dimensionality): the stability guarantee is exactly the same for any Hilbert space, regardless of dimension, as long as the radius of the output space $\mathcal{W}$ is bounded.
@@ 4.1 Results for a finite number of bags
The results above apply to the derandomized bagged version of $\mathcal{A}$, as constructed in Definition 5. In practice, of course, it is not feasible to calculate this output $\hat{w}$ since we would need to average over all possible bags $r$, so we would instead use the bagged estimator $\hat{w}$ constructed in Definition 4 for some finite number of bags $B$. Since we are simply taking a Monte Carlo approximation to the derandomized bagged estimator of Definition 5, we would expect to incur error on the order of $\frac{1}{\sqrt{B}}$. To formalize this, we first recall a generalization of the Azuma-Hoeffding inequality:
@@
Let $\hat{w}_{1},\ldots,\hat{w}_{B}\in\mathcal{H}$ be independent random variables in a Hilbert space $\mathcal{H}$ such that $\|\hat{w}_{b}-\mathbb{E}[\hat{w}_{b}]\|_{\mathcal{H}}\leq C$ almost surely for $b=1,\ldots,B$. Then for any $\delta>0$,
$\mathbb{P}\left\{\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}[\hat{w}_{b}])\bigg{\|}_{\mathcal{H}}\leq C\sqrt{\frac{1}{2B}\log\left(\frac{2e^{2}}{\delta}\right)}\right\}\geq 1-\delta.$
@@
$\mathbb{E}\left[\bigg{\|}\frac{1}{B}\sum_{b=1}^{B}(\hat{w}_{b}-\mathbb{E}[\hat{w}_{b}])\bigg{\|}_{\mathcal{H}}^{2}\right]\leq\frac{C^{2}e^{2}}{B}.$
@@ 5.1 Theoretical guarantee
Here $R$ is playing a role analogous to the radius, $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})$, in the Hilbert space setting. For instance, returning to our earlier example, if $\mathcal{W}=\Delta_{K-1}$ (and the norm is given by $\mathsf{d}_{\mathsf{TV}}$) then we can take $v_{j}$ to be the $j$th canonical basis vector. Then 2 holds, with this same value of $K$, and with $R=1$ and $\rho=0$.
@@ 5.2 Examples
Next, consider a continuous setting, where $\mathcal{A}$ returns a density on $[0,1]$, given by some function $\hat{w}:[0,1]\rightarrow[0,\infty)$ satisfying $\int_{t=0}^{1}\hat{w}(t)\;\textnormal{d}t=1$. In this case we have $\mathsf{d}_{\mathsf{TV}}(w,w^{\prime})=\frac{1}{2}\int_{t=0}^{1}|w(t)-w^{\prime}(t)|\;\textnormal{d}t$.
@@
$\mathcal{B}=L_{1}([0,1]):=\left\{f:[0,1]\to\mathbb{R}:\int_{t=0}^{1}|f(t)|\;\textnormal{d}t<\infty\right\},$
equipped with the norm $\|f\|=\frac{1}{2}\int_{t=0}^{1}|f(t)|\;\textnormal{d}t$ (note that this differs from the usual $\ell_{1}$ norm by the factor $\frac{1}{2}$, in order to agree with total variation distance). Suppose that our algorithm returns densities that are further constrained to be $L$-Lipschitz, so that we have
@@
As for Problem Setting 2(a), we again begin by asking whether the Hilbert space result of Theorem 6 might already be sufficient in this setting. Indeed, while $\mathcal{B}=L_{1}([0,1])$ is not a Hilbert space, the Lipschitz constraint on $\mathcal{W}$ means that $\int_{t=0}^{1}|w(t)|^{2}\;\textnormal{d}t<\infty$ for all $w\in\mathcal{W}$, meaning that $\mathcal{W}$ can be viewed as a subset of the Hilbert space $L_{2}([0,1])$, equipped with norm $\|f\|_{\mathcal{H}}=\|f\|_{L_{2}}=(\int_{t=0}^{1}|f(t)|^{2}\;\textnormal{d}t)^{1/2}$. A straightforward calculation verifies that $\mathsf{rad}_{\mathcal{H}}(\mathcal{W})\propto L^{1/4}$, meaning that by Theorem 6, derandomized bagging satisfies mean-square stability with respect to $\|\cdot\|_{L_{2}}$ for $\beta^{2}\propto L^{1/2}/n$. Since for this function space we have $\|\cdot\|_{L_{1}}\leq\|\cdot\|_{L_{2}}$, this immediately implies that mean-square stability with respect to $\mathsf{d}_{\mathsf{TV}}$ also holds with parameter $\beta^{2}\propto L^{1/2}/n$. However, if the Lipschitz constant $L$ is large, this may not be a satisfying result.
@@
Stability guarantee for $\mathsf{d}_{\mathsf{TV}}$
via Theorem 6
(Hilbert space)
Stability guarantee for $\mathsf{d}_{\mathsf{TV}}$
via Theorem 10
(Banach space)
@@ 6.1 Data and Methods
with $n=500$ and $d=40$. We then set $Y_{i}=\frac{\tilde{Y}_{i}-\min_{j}\tilde{Y}_{j}}{\max_{j}\tilde{Y}_{j}-\min_{j}\tilde{Y}_{j}}$ so that each response $Y_{i}$ is in the unit interval $[0,1]$. Note that the algorithm has access to the observed data $\mathcal{D}=(X_{i},Y_{i})_{i\in[n]}$, i.e., $\alpha_{i}$ and $\gamma_{i}$ are latent variables used only to generate the data $\mathcal{D}$. We apply sklearn.tree.DecisionTreeRegressor to train the regression trees, setting max_depth=50 and leaving all other parameters at their default values. We evaluate the stability of the learned function in the $L_{2}$ norm $\|\hat{w}\|_{L_{2}}=(\int_{[0,1]^{d}}\hat{w}(x)^{2}\textnormal{d}x)^{1/2}$.
@@
For this experiment, we reproduce Abadie et al. (2015)’s SCM analysis on the 1990 reunification of Germany, as discussed above in Section 1.1 (see Problem Setting 3). The data contains $n=16$ countries (aside from Germany), with $6$ predictors measured for each country (pre-treatment averages of various macroeconomic factors such as inflation and trade openness—see Table 1 of Abadie et al. (2015) for more details). SCM then estimates a weighted combination of these $n$ countries that represents a ‘synthetic West Germany’, meaning that the output of the algorithm is a vector $\hat{w}\in\Delta_{n-1}$. When we run SCM on a subset of the $n=16$ countries, we assign a weight of 00 to each omitted country. We run SCM using the Python package pysyncon and make extensive use of their reproduction of the original study Abadie et al. (2015). We evaluate the stability of these weights in the Euclidean norm $\|\cdot\|_{2}$.
@@
Figure 1: Simulation results comparing the stability of subbagging to that of the corresponding base algorithm under the given norm. Left: Histogram of leave-one-out perturbations. Right: for each $\varepsilon$, the smallest $\delta$ such that the algorithm has tail stability $(\varepsilon,\delta)$ in the sense of Definition 2. Higher curves thus represent greater instability. In the right panels we also plot the stability guarantee of Theorem 6 for Experiments 1–3 and Theorem 10 for Experiment 4 (the guarantees for derandomized bagging). For all four experiments, $m=n/2$ and $B=10,000$.